OpenStudy (anonymous):

Twenty pounds of a salt-water solution contains fifteen percent salt. How much water must be added to weaken the solution to twelve percent salt? i got 12 lb.

OpenStudy (lgbasallote):

water is 0% salt...did you use that?

OpenStudy (anonymous):

3 lb?.

OpenStudy (anonymous):

um i'm not sure..

OpenStudy (lgbasallote):

i'll see what @Shane_B has to say...

OpenStudy (anonymous):

There are 20 x 0.15 = 3 lbs of salt before dilution, and of course 3 lbs afterward, too. So the question becomes 3/(20 + x) = 0.12, solve for x.

OpenStudy (shane_b):

That's right...had to think about it (it's 5:30am here :)

OpenStudy (anonymous):

so its 12lb.

OpenStudy (anonymous):

No, 5. 3/(20 + 5) = 3/25 = 0.12.